### v4

# Coupling without Communication and Drafter-Invariant Speculative Decoding

Abstract
Suppose Alice has a distribution $\mathcal{P}$ and Bob has a distribution $\mathcal{Q}$. Alice wants to draw a sample $a\sim\mathcal{P}$ and Bob a sample $b\sim\mathcal{Q}$ such that $a=b$ with as high of probability as possible. It is well-known that, by sampling from an optimal coupling between the distributions, Alice and Bob can achieve $\Pr[a=b]=1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$, where $D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$ is the total variation distance between $\mathcal{P}$ and $\mathcal{Q}$.
What if Alice and Bob must solve this same problem without communicating at all? Perhaps surprisingly, with access to public randomness, they can still achieve $\Pr[a=b]\geq\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}\geq 1-2D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$ using a simple protocol based on the Weighted MinHash algorithm. This bound was shown to be optimal in the worst-case by Bavarian, Ghazi, Haramaty, Kamath, Rivest, and Sudan [ToC 2020].
In this work, we revisit the “communication-free coupling” problem. We provide a simpler proof of the optimality result from [Bavarian et al., 2020]. Moreover we show that, while the worst-case success probability of Weighted MinHash cannot be improved, an equally simple protocol based on Gumbel sampling offers a Pareto improvement: for every pair of distributions $\mathcal{P}$ and $\mathcal{Q}$, Gumbel sampling achieves an equal or higher value of $\Pr[a=b]$ than Weighted MinHash.
Importantly, this improvement translates to practice. We demonstrate an application of communication-free coupling to speculative decoding, a recent method for accelerating autoregressive large language models [Leviathan, Kalman, Matias, ICML 2023]. We show that communication-free protocols can be used to contruct Drafter-Invariant Speculative Decoding schemes, which have the desirable property that their output is fixed given a fixed random seed, regardless of what drafter is used for speculation. In experiments on a language generation task, Gumbel sampling outperforms Weighted MinHash. Code is available at https://github.com/majid-daliri/DISD.
Finally, we study the coupling problem in the setting where communication is bounded, rather than completely eliminated. We describe a protocol that uses just $O(\log(n/\epsilon))$ bits of communication to achieve $\Pr[a=b]=1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})-\epsilon$, i.e. to essentially match optimal coupling.
We thank Cheuk Ting Li for bringing several related prior works to our attention after we initially published this paper. Angel and Spinka (2021) prove the same result that we do on the Gumbel sampling method, showing that it offers a Pareto improvement over Weighted MinHash. Moreover, in addition to our work and the cited work by Bavarian et al. (2020), Li and Anantharam (2019, 2021) also prove that $\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}$ is optimal in the worst-case for communication-free coupling.
1 Introduction
The goal of this paper is to revist a “communication-free coupling” problem that arises repeatedly in computer science, and was studied explicitly by Bavarian et al. (2020) under the name correlated sampling. Suppose we have two cooperating parties, Alice and Bob. Alice has a distribution $\mathcal{P}$ and Bob has a distribution $\mathcal{Q}$, which, for simplicity, we assume are discrete distributions over the set $\{1,\ldots,n\}$. The goal is for Alice to sample $a\in\{1,\ldots,n\}$ according to $\mathcal{P}$ and Bob to sample $b\in\{1,\ldots,n\}$ according to $\mathcal{Q}$, in such a way that we maximize the probability that $a=b$. We are interested in how well Alice and Bob can solve this problem without communicating with each other. We do allow them access to a source of shared random numbers.
If communication was allowed, it is well known that Alice and Bob can achieve:
$\Pr[a=b]=1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q}),$ (1)
where $D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$ is the total variation distance between $\mathcal{P}$ and $\mathcal{Q}$ (Wu, 2020, Section 4.3). To do so, they construct an optimal coupling between the distributions, i.e., a distribution $\mathcal{C}$ over $\{1,\ldots,n\}\times\{1,\ldots,n\}$ with marginals $\mathcal{P}$ and $\mathcal{Q}$ that maximizes $\Pr_{(a,b)\sim\mathcal{C}}[a=b]$. In fact, the total variation distance is defined as $\max_{\text{couplings }\mathcal{C}}\Pr_{(a,b)\sim\mathcal{C}}[a\neq b]$, so (1) is optimal by definition, with or without communication.
While an optimal coupling between $\mathcal{P}$ and $\mathcal{Q}$ has a simple closed form (see Section 2 for details), sampling from the coupling requires knowledge of both $\mathcal{P}$ and $\mathcal{Q}$. I.e., sampling from an optimal coupling requires communication between Alice and Bob. However, it turns out that it is possible to sample from a near-optimal coupling without any communication at all. In particular, the starting point of our work is a straightforward observation from Bavarian et al. (2020), which is implicit in prior work (Manasse et al., 2010):
Fact 1.
There is a communication-free protocol (Protocol 2) by which Alice samples $a\sim\mathcal{P}$ and Bob $b\sim\mathcal{Q}$ such that:
$\Pr[a=b]\geq\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}.$ (2)
Since $0\leq D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})\leq 1$, (2) is always larger than $1-2D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$. We conclude that it is possible to almost match the optimum in (1) without communication. 1 follows from an analysis of the popular “Weighted MinHash” method (Kleinberg and Tardos, 2002; Holenstein, 2009; Manasse et al., 2010; Ioffe, 2010; Shrivastava, 2016; Christiani, 2020). We provide a sharp analysis of the method (with an exact expression for $\Pr[a=b]$) in Section 3, 5.
1.1 Motivation
The main question we address in this work is if it is possible to improve even further on the excellent performance of the existing Weighted MinHash method for communication-free coupling. To cut to the point: it is. We show that a simple method based on the “Gumbel Max-Trick” provides a Pareto improvement over Weighted MinHash. Before discussing this result in detail, however, we provide some motivation for studying communication-free coupling in such depth. It turns out that this problem has a direct application to a technique called speculative decoding that has gained traction for its ability to accelerate inference in autoregressive language models (Leviathan et al., 2023; Chen et al., 2023; Sun et al., 2023), one of the most important algorithmic challenges in modern machine learning (Hoffmann et al., 2022; Zhou et al., 2024).
Details of speculative decoding are given in Section 5. In short, the method speeds up language models by using a small and inexpensive neural network to draft (i.e., speculate) the tokens that will be generated by a larger state-of-the-art network. Tighter coupling between the token distributions of the two networks leads to higher chance of draft tokens matching the larger model tokens, which leads to accelerated inference.
However, speculative decoding suffers from one major and subtle drawback: if the small “drafter” neural network changes (i.e., because of an update to improve the model), the use of an optimal coupling means that tokens generated by the state-of-the-art neural network will change. This can be potentially problematic in applications of autoregressive language models, where the expected and desired behavior is that the output of a model is fixed given a fixed random seed. This behavior allows for researchers and other users to reliably reproduce results and leads to easier unit testing and debugging.
As we will demonstrate in Section 5, communication-free coupling allows for the implementation of a Drafter-Invariant Speculative Decoding method that avoids this issue. In particular, the output of the state-of-the-art model is completely independent of what drafter network is used to accelerate inference: it is fixed given a fixed random seed. Intuitively this follows from the fact that in any communication-free coupling protocol, Alice and Bob’s samples are necessarily independent when conditioned on their shared randomness.
1.2 Our Contributions
Motivated by this application, our work seeks to understand if it is possible to improve on the existing Weighted MinHash method for communication-free coupling. Increasing $\Pr[a=b]$ from $\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}$ closer to the optimal $1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$ could mean up to a 2x reduction in wrong guesses during speculative decoding, which would translate to a potential 2x reduction in computational depth during language generation.
Unfortunately, improving on the bound in 1 is not possible in the worst-case. As shown in an elegant proof by Bavarian et al. (2020), no protocol can do better than $\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}$ in the worst-case. We restate a version [^1] The result in Bavarian et al. (2020) is slightly stronger than our Theorem 2 in that it allows for any total variation distance, not just distances of the form $1/d$ for integer $d$. We assume distance of this form to simplify our argument. of this result below and provide an even more compact proof in Section 3.2.
Theorem 2.
Consider any protocol that takes as input a distribution and source of public randomness. For any positive integer $d$, there are distributions $\mathcal{P}$ and $\mathcal{Q}$ with total variation distance $D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})=1/d$ such that, if Alice runs the protocol to sample $a\sim\mathcal{P}$ and Bob runs the protocol to sample $b\sim\mathcal{Q}$, then:
$\Pr[a=b]\leq\frac{1-1/d}{1+1/d}.$
Importantly, while Theorem 2 rules out an improvement on Weighted MinHash for all distribution pairs, it does not mean that the method is optimal. Our main contribution is to demonstrate that another extremely simple protocol offers a Pareto improvement over Weighted MinHash: for any $\mathcal{P},\mathcal{Q}$ it performs at least as well, but sometimes performs better (often significantly so in practice). Our alternative approach is based on the so-called “Gumbel Max-Trick” or “Gumbel sampling”. The method can be described in one sentence: using shared random variables $u_{1},\ldots,u_{n}$ drawn uniformly from $[0,1]$, Alice returns $a=\operatorname{argmin}_{i\in\{1,\ldots,n\}}\frac{-\ln(u_{i})}{p_{i}}$  [^2] Often $\operatorname{argmin}_{i\in\{1,\ldots,n\}}\frac{-\ln(u_{i})}{p_{i}}$ is computed as $\operatorname{argmax}_{i\in\{1,\ldots,n\}}\ln p_{i}-\ln(-\ln(u_{i}))$. If $u$ is a uniform random variable in $[0,1]$, then $\ln(-\ln(u))$ follows a Gumbel distribution (Gumbel, 1935) and hence the name. and Bob returns $b=\operatorname{argmin}_{i\in\{1,\ldots,n\}}\frac{-\ln(u_{i})}{q_{i}}$, where $p_{i}$ and $q_{i}$ denote the respective probabilities that Alice and Bob’s distributions, $\mathcal{P}$ and $\mathcal{Q}$, assign to item $i$. Intuitively, $a$ and $b$ will be partially coupled because, if some $u_{i}$ happens to be small, we are more likely to set both $a=i$ and $b=i$.
Gumbel sampling has been widely studied in theoretical computer science under different names, including PPSWOR and bottom-k sampling (Cohen, 1997; Rosén, 1997; Cohen, 2023). Moreover, it is already widely used in machine learning for uncoupled sampling from discrete distributions, including for auto-regressive language generation (Kool et al., 2019). So, using the method for speculative decoding would require essentially no code changes (Maddison et al., 2014; Huijben et al., 2023). We prove the following bound on the performances of Gumbel sampling in Section 3:
Theorem 3.
For distributions $\mathcal{P}$ and $\mathcal{Q}$, let $C_{\text{WMH}}(\mathcal{P},\mathcal{Q})$ be the probability that samples $a\sim\mathcal{P}$ and $b\sim\mathcal{Q}$ drawn using the Weighted MinHash method (Protocol 2) satisfy $a=b$. The communication-free Gumbel coupling method (Protocol 3) samples $a\sim\mathcal{P}$ and $b\sim\mathcal{Q}$ such that:
$\Pr[a=b]=\sum_{\begin{subarray}{c}j\in\{1,\ldots,n\}\\ \min(p_{j},q_{j})>0\end{subarray}}\frac{1}{\sum_{i=1}^{n}\max(p_{i}/p_{j},q_{i}/q_{j})}\geq C_{\text{WMH}}(\mathcal{P},\mathcal{Q})\geq\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}.$
It is not hard to find distributions for which the first inequality above is strict, i.e., for which Gumbel sampling strictly improves on Weighted MinHash. In fact, as discussed in Section 3, this will always be the case under the mild condition that $p_{i}\neq q_{i}$ for at least three values of $i$. As an example, consider distributions $\mathcal{P}=\{{1}/{2},{1}/{2},0\}$ and $\mathcal{Q}=\{{1}/{3},{1}/{3},{1}/{3}\}$ over $\{1,2,3\}$. The best possible collision probability for these distributions is $1-\mathcal{D}_{TV}(\mathcal{P},\mathcal{Q})={2}/{3}$ and $\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}=1/2$. In this case, Gumbel sampling actually obtains the optimal collision probability of 2/3{2}/{3}, while Weighted MinHash obtains probability $7/12$, closer to the worst-case bound for communication-free methods. A similar pattern emerges in our experiments with real-word next-token distributions that arise in language generation Section 5. Gumbel sampling often nearly matches the optimal coupling, obtaining collision probability close to $1-\mathcal{D}_{TV}(\mathcal{P},\mathcal{Q})$, while Weighted MinHash performs somewhat worse.
It is interesting to ask if additional improvement is possible. Is there another protocol that offers a Pareto improvement over Gumbel sampling, or is the approach Pareto optimal? I.e., does improvement for a pair of distributions $(\mathcal{P},\mathcal{Q})$ necessarily require worse performance for another pair?
Considering the worst-case bound of Theorem 2, it is also natural to ask what is possible if we restrict, but do not entirely eliminate communication. How many bits of communication between Alice and Bob are required to ensure that $\Pr[a=b]=1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})-\epsilon$, i.e., to nearly match what is possible with optimal coupling? A baseline is to simply discretize $\mathcal{P}$ or $\mathcal{Q}$, communicate the entire distribution, and compute a near optimal coupling. This takes $O(n\log(n/\epsilon))$ bits of communication (see Section 6 for details). We show that it is possible to get by with much less. In Section 6 we prove:
Theorem 4.
There is a protocol that, for any $\epsilon\in(0,1)$, requires a constant number of rounds and $O(\log(n/\epsilon))$ bits of communication, in expectation, between Alice and Bob to produce samples $a\sim\mathcal{P}$, $b\sim\mathcal{Q}$ such that:
$\Pr\left[a=b\right]\geq 1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})-\epsilon.$
While we are not aware of immediate applications of Theorem 4, it provides an interesting point of comparison to both the full-communcation and no-communcation settings.
Finally, beyond our theoretical results, we implement our Gumbel sampling method and evaluate its performance in the “drafter-invariant speculative decoding” task in Section 5. We show that the method nearly matches standard speculative decoding in terms of speculation accuracy, but with the added benefit that the neural network output is completely invariant to the drafter network utilized.
1.3 Related Work
Our work is closely related to results on weighted coordinated sampling methods, which generalize well-known techniques like MinHash and the $k$-minimum values sketch for unweighted coordinated sampling (Broder, 1997; Broder et al., 1998; Cohen and Kaplan, 2007; Beyer et al., 2007; Li et al., 2006). The goal in weighted coordinated sampling is similar to ours: Alice and Bob hold vectors, $A$ and $B$, and seek to independently produce a subsample of entries from their vector so that 1) indices corresponding to larger entries in the vectors are sampled with higher probability and 2) Alice and Bob return many of the same indices. The aim is to use Alice and Bob’s subsamples to estimate functions involving interactions between corresponding entries in $A$ and $B$, like the inner product $\langle A,B\rangle=\sum_{i=1}^{n}A_{i}B_{i}$ (Li, 2017; Bessa et al., 2023; Daliri et al., 2024a). To do so effectively, it is critical that we have access to pairs $(A_{i},B_{i})$ with the same index $i$.
Weighted coordinated sampling methods include Weighted MinHash (also referred to as “consistent weighted sampling”) (Manasse et al., 2010; Ioffe, 2010; Haeupler et al., 2014; Wu et al., 2020), Gumbel sampling (Cohen, 2015, 2023), threshold sampling (Flajolet, 1990; Duffield et al., 2005), priority sampling (Duffield et al., 2004; Daliri et al., 2024b), and more (Estan and Naughton, 2006). While all of these methods intuitively seek to generate samples from $A$ and $B$ that contain the same indices with high probability, that is typically not the final goal: the methods are analyzed for specific downstream applications. As such we are unaware of any prior work besides that of Bavarian et al. (2020) which specifically obtains bounds for our “communication-free coupling” problem.
We also note that several coordinated sampling methods cannot be applied in our context. For example, priority sampling does not ensure that entries from $A$ and $B$ are sampled with probability exactly proportional to specified probabilities (we require samples to truly be drawn from $\mathcal{P}$ and $\mathcal{Q}$), and threshold sampling does not return a fixed number of samples (we always need exactly one sample). Of the major methods, this leaves Weighted MinHash and Gumbel sampling, which are the two methods we address in this paper.
2 Preliminaries
Notation. Throughout the paper, we assume two parties, Alice and Bob, who hold discrete distributions $\mathcal{P}$ and $\mathcal{Q}$ over the set $\{1,\ldots,n\}$. We use $p_{i}$ to denote the probability that $x$ drawn from $\mathcal{P}$ equals $i$ and $q_{i}$ to denote the probability that $x$ drawn from $\mathcal{Q}$ equals $i$. We will sometimes write $\mathcal{P}=\{p_{1},\ldots,p_{n}\}$ and $\mathcal{Q}=\{q_{1},\ldots,q_{n}\}$ since the list of probabilities completely specifies the distributions.
Total Variation Distance. The total variation distance, $D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$, between discrete distributions equals:
$D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})=\frac{1}{2}\sum_{i=1}^{n}|p_{i}-q_{i}|=\sum_{i=1}^{n}\max(0,p_{i}-q_{i})=\sum_{i=1}^{n}\max(0,q_{i}-p_{i}).$ (3)
Throughout our proofs, we will use two elementary inequalities that follow from (3):
$1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$ $=\sum_{i=1}^{n}p_{i}-\max(0,p_{i}-q_{i})=\sum_{i=1}^{n}\min(p_{i},q_{i})$ (4)
$1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$ $=\sum_{i=1}^{n}p_{i}+\max(0,q_{i}-p_{i})=\sum_{i=1}^{n}\max(p_{i},q_{i})$ (5)
Coupling with Communication. As discussed in Section 1, if they are allowed to communicate, Alice and Bob can easily sample from $a\sim\mathcal{P}$ and $b\sim\mathcal{Q}$ in such a way that $\Pr[a=b]=1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$. Concretely, they can execute the following standard protocol:
Protocol 1 Coupling with Communication (Leviathan et al., 2023; Chen et al., 2023)
Protocol for Alice (who has probability vector $\mathcal{P}=[p_{1},\ldots,p_{n}]$):
1:Sample $a\sim\mathcal{P}$. Communicate $a$ and the distribution $\mathcal{P}$ to Bob.
Protocol for Bob (who has probability vector $\mathcal{Q}=[q_{1},\ldots,q_{n}]$):
1:Await for $(a,\mathcal{P})$ from Alice. With probability $\min(1,q_{a}/p_{a})$ set $b=a$.
2:Otherwise, sample $b\sim\mathcal{Q}^{\prime}$, where $\mathcal{Q}^{\prime}=\{q_{1}^{\prime},\ldots,q_{n}^{\prime}\}$ is a distribution with $q_{i}^{\prime}=\frac{\max(0,q_{i}-p_{i})}{\sum_{j=1}^{n}\max(0,q_{j}-p_{j})}$.
Clearly Protocol 1 ensures that $a\sim\mathcal{P}$. It is easily checked that it also ensures that $b\sim\mathcal{Q}$. In particular, for any $i$ where $q_{i}\leq p_{i}$, $\max(0,q_{i}-p_{i})=0$, so we can only set $b=i$ in Step 1 of the protocol. We do so with probability $\Pr[a=i]\cdot q_{i}/p_{i}=q_{i}$, as desired. If $q_{i}\geq p_{i}$, then with probability $p_{i}$, we set $b=i$ in Step 1 of the protocol (exactly when $a$ is set to $i$). There is also some chance we set $b=i$ in Step 2 of the protocol, which we execute with probability $\sum_{j=1}^{n}p_{j}\cdot(1-\min(1,q_{j}/p_{j}))=\sum_{j=1}^{n}\max(0,p_{j}-q_{j})=\sum_{j=1}^{n}\max(0,q_{j}-p_{j})$. So, the overall probability we set $b=i$ in the second step is $q_{i}^{\prime}\cdot\sum_{j=1}^{n}\max(0,q_{j}-p_{j})=q_{i}-p_{i}$. Thus, the total probability we set $b=i$ (in either step) is $q_{i}-p_{i}+p_{i}=q_{i}$, as desired.
Finally, we can see that $\Pr[a=b]=\sum_{j=1}^{n}p_{j}\cdot\min(1,q_{j}/p_{j})=\sum_{j=1}^{n}\min(p_{j},q_{j})=1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$ via (4).
3 An Existing Communication-free Protocol
As discussed, a communication-free protocol based on the Weighted MinHash algorithm can nearly match the performance of Protocol 1 (which requires communication, since Bob needs to know Alice’s distribution $\mathcal{P}$). A number of papers analyze Weighted MinHash, including Bavarian et al. (2020). We provide a self-contained analysis below for completeness, and to obtain a sharp bound (i.e., an exact expression for $\Pr[a=b]$).
3.1 Weighted MinHash
Technically speaking, there are many different ways to implement Weighted MinHash, which might be more accurately described as a family of closely-related sampling method. See (Wu et al., 2020) for an overview. We analyze a particularly simple implementation suggested by a number of authors (Kleinberg and Tardos, 2002; Holenstein, 2009; Shrivastava, 2016). It is detailed in Protocol 2 and illustrated in Figure 1.
Protocol 2 Weighted MinHash Coupling
Fix public random numbers $u_{1},u_{2},u_{3}\ldots$ drawn uniformly from the interval $[0,n]$.
Protocol for Alice (who has probability vector $\mathcal{P}=[p_{1},\ldots,p_{n}]$):
1:for $k=1,2,\ldots$ do
2:  if $u_{k}\in[j-1,j-1+p_{j}]$ for some $j\in\{1,\ldots,n\}$ then return $a=j$.
Protocol for Bob (who has probability vector $\mathcal{Q}=[q_{1},\ldots,q_{n}]$):
1:for $k=1,2,\ldots$ do
2:  if $u_{k}\in[j-1,j-1+q_{j}]$ for some $j\in\{1,\ldots,n\}$ then return $b=j$.
Figure 1: The Weighted MinHash method, described in Protocol 2, selects a sample from a distributions on $n$ items with probabilities $p_{1},\ldots,p_{n}$ by drawing a sequence of numbers $u_{1},u_{2},\ldots,$ uniformly at random from the interval $[0,n]$. The first time one of these numbers lands in a subinterval $j-1,j-1+p_{j}]$ for any $j\in 1,\ldots,n$ , the protocol returns $j$. For example, in the illustration above, Alice returns 22 (since $u_{2}$ falls within $1,1+p_{2}]$) and Bob returns 44 (since $u_{5}$ falls within $3,3+p_{4}]$)
Claim 5.
The communication-free Weighted MinHash coupling method (Protocol 2) generates $a\sim\mathcal{P}$, $b\sim\mathcal{Q}$ such that:
$\Pr[a=b]=\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})+\sum_{i=1}^{n}|p_{i}-q_{i}|\min(p_{i},q_{i})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}\geq\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}.$
Proof.
As illustrated in Figure 1, to run Protocol 2, Alice draws values uniformly at random from the interval $[0,n]$. The protocol terminates and returns $j$ as soon as a value lands in any subinterval $j-1,j-1+p_{j}]$, which are shaded in the top row of Figure 1. Since these subintervals have lengths $p_{1},\ldots,p_{n}$, the probability that Alice returns $j$ is exactly $\frac{p_{j}}{\sum_{i=1}^{n}p_{i}}=p_{j}$. The proof that Bob returns $j$ with probability $q_{j}$ is identical.
We turn our attention to the expression for $\Pr[a=b]$. Let $u_{k}$ denote be the first uniform random number for which either Alice or Bob terminates, i.e., the first $k$ for which $u_{k}\in[j-1,j-1+\max(p_{j},q_{j})]$ for some $j$. We split the analysis into two cases:
Case 1.
Both Alice and Bob terminate at step $k$, i.e. $u_{k}\in[j-1,j-1+\min(p_{j},q_{j})]$.
Case 2.
Only one of Alice and Bob terminate at step $k$, i.e. $u_{k}\in[j-1+\min(p_{j},q_{j}),j-1+\max(p_{j},q_{j})]$.
In the first case, they always return the same index, $j=\lceil u_{k}\rceil$, so we have that $a=b$. Additionally, we can see that, conditioned on the event that $k$ is the first index for which $u_{k}\in[j-1,j-1+\max(p_{j},q_{j})]$ for some $j$ (i.e., that $k$ is the first step that someone terminates at), the first case happens with probability:
$\frac{\sum_{i=1}^{n}\min(p_{i},q_{i})}{\sum_{i=1}^{n}\max(p_{i},q_{i})}.$ (6)
Next consider the second case. In this case, only one of Alice and Bob terminates at iteration $k$. The other terminates at iteration $k^{\prime}>k$, and it may be that $a\neq b$. To analyze the collision probability, we further break down the case, considering only when $u_{k}\in[j-1+\min(p_{j},q_{j}),j-1+\max(p_{j},q_{j})]$ for a particular value of $j$. We can see that, conditioned on the fact that one of the parties terminates at step $k$,
$\Pr\left[u_{k}\in[j-1+\min(p_{j},q_{j}),j-1+\max(p_{j},q_{j})]\right]=\frac{|p_{j}-q_{j}|}{\sum_{i=1}^{n}\max(p_{i},q_{i})}.$
If $u_{k}$ does land in the interval $j-1+\min(p_{j},q_{j}),j-1+\max(p_{j},q_{j})]$, Alice is the party who terminates and return $j$ whenever $p_{j}>q_{j}$. Otherwise, if $q_{j}>p_{j}$, Bob terminates and returns $j$. So, the question becomes, with what probability does the other party also returns $j$ on a future iteration? Future iterations involve random variables completely independent from $u_{k}$, so the probability is simply $\min(p_{j},q_{j})$.
Putting together both cases, we have that:
$\Pr[a=b]=\frac{\sum_{i=1}^{n}\min(p_{i},q_{i})}{\sum_{i=1}^{n}\max(p_{i},q_{i})}\cdot 1+\sum_{j=1}^{n}\frac{|p_{j}-q_{j}|}{\sum_{i=1}^{n}\max(p_{i},q_{i})}\cdot\min(p_{j},q_{j}).$
The final result follows from rearranging and simplifying by applying Equations 4 and 5. ∎
3.2 Optimality of Communication-free Protocols
As discussed in Section 1, when parameterizing performance by total variance distance, it can be proven that the bound of 5 for Weighted MinHash is optimal for worst-case distributions. Bavarian et al. (2020) do so via a reduction from an easier “constrained agreement problem”. We provide an even simpler proof here, first restating the formal optimality claim from Section 1:
See 2
Proof.
As a warm up, we first consider a case involving distributions with total variation distance $1/2$, i.e., $d=2$. We will show that, without communication, it is not possible to achieve $\Pr[a=b]\geq\frac{1-1/2}{1+1/2}=1/3$. Specifically consider the following distributions $\mathcal{P},\mathcal{Q}$ and $\mathcal{R}$, which are each supported on three elements:
$\begin{bmatrix}p_{1}\\ p_{2}\\ p_{3}\end{bmatrix}$ $=\begin{bmatrix}1/2\\ 1/2\\ 0\end{bmatrix}$ $\begin{bmatrix}q_{1}\\ q_{2}\\ q_{3}\end{bmatrix}$ $=\begin{bmatrix}1/2\\ 0\\ 1/2\end{bmatrix}$ $\begin{bmatrix}r_{1}\\ r_{2}\\ r_{3}\end{bmatrix}$ $=\begin{bmatrix}0\\ 1/2\\ 1/2\end{bmatrix}$.
Now, suppose a protocol is used to sample $a\sim\mathcal{P},b\sim\mathcal{Q}$ and $c\sim\mathcal{R}$. For any $i\in\{1,2,3\}$, it must be that one of $a,b$ or $c$ does not equal $i$, since at least one of the distributions has 0 mass on item $i$. It follows that only one of the equalities $a=b$, $a=c$, and $b=c$ holds at once. So, we have that for any protocol, $\min(\Pr[a=b],\Pr[a=c],\Pr[b=c])\leq 1/3$. I.e., for two of the distributions, the protocol gives collision probability $\leq 1/3$. The result then follows from simply verifying that that $D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})=D_{\mathrm{TV}}(\mathcal{P},\mathcal{R})=D_{\mathrm{TV}}(\mathcal{Q},\mathcal{R})=1/2$:
We can generalize this result by having $d+1$ distributions $\mathcal{P}_{1},\ldots,\mathcal{P}_{d+1}$ over $1,\ldots,d+1$. Specifically, let the $i^{\text{th}}$ distribution, $\mathcal{P}_{i}$, be uniform over all $j\neq i$. So, $x\sim\mathcal{P}_{i}$ has $1/d$ probability of equaling any $j\neq i$. It can be checked that $D_{\mathrm{TV}}(\mathcal{P}_{i},\mathcal{P}_{j})=1/d$ for all $i,j$.
Now, consider a protocol that samples $x_{1}\sim\mathcal{P}_{1},\ldots,x_{d+1}\sim\mathcal{P}_{d+1}$. Consider all $d+1\choose 2$} equalities of the form $\{x_{i}=x_{j},i\neq j\}$. No matter what the outcome of $x_{1},\ldots,x_{d+1}$, we claim that at least $d$ of these inequalities will not be satisfied. In particular, as in the argument above, it cannot be that $x_{1},\ldots,x_{d}$ all have the same value $v$ since one of $\mathcal{P}_{1},\ldots,\mathcal{P}_{d+1}$ has zero mass for any possible $v$. As such, at least two items in $x_{1},\ldots,x_{d+1}$ must be different from each other. Suppose without loss of generality that $x_{1}\neq x_{2}$.
First, the equality $x_{1}=x_{2}$ is immediately not satisfied. Then for all $j\neq 1,2$, it must be that either $x_{j}\neq x_{1}$ or $x_{j}\neq x_{2}$, which is d+12=d1d+1-2=d-1 additional inequalities not satisfied. So, no matter what the outcome of the protocol, the ratio of equalities of the form $\{x_{i}=x_{j},i\neq j\}$ that are satisfied is at most:
$1-\frac{d}{{d+1\choose 2}}=1-\frac{2}{d+1}.$
It follows that there must be some pair $i,j$ such that:
$\Pr[x_{i}=x_{j}]\leq 1-\frac{2}{d+1}=\frac{1-1/d}{1+1/d}.$
4 Better Coupling via Gumbel Sampling
Despite the limitation of LABEL:{clm:ub}, in this section we show that it is nevertheless possible to strictly improve on the Weighted MinHash method from the previous section. To do so, we analyze a method based on Gumbel sampling, an existing protocol that is widely used for sampling from discrete distributions, even when coupling is not necessary. Under various names like PPSWOR, order sampling, and bottom-k sampling, the method has found widespread application across randomized algorithms (Rosén, 1997; Cohen, 2023; Efraimidis and Spirakis, 2006). We give pseudocode in Protocol 3.
Protocol 3 Gumbel Coupling
Fix public random numbers $u_{1},\ldots,u_{n}$ drawn uniformly from the interval $[0,1]$.
Protocol for Alice (who has probability vector $\mathcal{P}=[p_{1},\ldots,p_{n}]$):
1:Return $a=\operatorname{argmin}_{i\in\{1,\ldots,n\}}\frac{-\ln(u_{i})}{p_{i}}$.
Protocol for Bob (who has probability vector $\mathcal{Q}=[q_{1},\ldots,q_{n}]$):
1:Return $b=\operatorname{argmin}_{i\in\{1,\ldots,n\}}\frac{-\ln(u_{i})}{q_{i}}$.
Our main result on Gumbel sampling is Theorem 3, which we restate below: See 3
As discussed in Section 1, it is not hard to find distributions for which the inequalities above are strict. In fact, Gumbel sampling often obtains a collision probability to the optimal $1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$. Understanding this strong practical performance is an interesting direction for future work. It may be be that the bound of Theorem 3 tends to be much stronger for the sort of “quickly decaying” distributions common in applications. As an extreme example, suppose that $n=2$, or equivalently, $\mathcal{P}$ and $\mathcal{Q}$ only place mass on two items. Then, as above, Gumbel sampling has exactly the same collision probability as that of the optimal coupling. To see this, without loss of generality, let $p_{1}\leq q_{1}$, then the result of Theorem 3 simplifies as:
$\Pr[a=b]=\frac{1}{1+\max\left(\frac{p_{1}}{p_{2}},\frac{q_{1}}{q_{2}}\right)}+\frac{1}{1+\max\left(\frac{p_{2}}{p_{1}},\frac{q_{2}}{q_{1}}\right)}=\frac{1}{1+\frac{q_{1}}{q_{2}}}+\frac{1}{1+\frac{p_{2}}{p_{1}}}=q_{2}+p_{1}=1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q}).$
We prove Theorem 3 in two parts, first showing that $\Pr[a=b]\geq\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}$. We then prove that $\Pr[a=b]\geq C_{\text{WMH}}$ which is more difficult. To do so, we require some standard preliminaries about exponential random variables. For a reference, see, e.g. Cohen (2023).
Fact 6 (Properties of Exponential Random Variables).
A random variable $X\sim\operatorname{Exp}(\lambda)$ is exponential with parameter $\lambda$ if $X=-\ln(u)/\lambda$ for a random variable $u$ drawn uniformly from $[0,1]$. Suppose we have independent exponential random variables $X_{1}\sim\operatorname{Exp}(\lambda_{1})$ and $X_{2}\sim\operatorname{Exp}(\lambda_{2})$. Then we have:
  • $\Pr[X_{1}<X_{2}]=\frac{\lambda_{1}}{\lambda_{1}+\lambda_{2}}$.
  • $\min(X_{1},X_{2})$ is an exponential random variable with parameter $\lambda_{1}+\lambda_{2}$.
4.1 Main Analysis
Proof of Theorem 3.
First note that:
$\Pr[a=b]=\sum_{j=1}^{n}\Pr[a=j\text{ and }b=j].$
So, we focus on analyzing $\Pr[a=j\text{ and }b=j]$ for a fixed $j$. Note that the probability is 0 is either $p_{j}=0$ or $q_{j}=0$, so we assume below that $\min(p_{j},q_{j})>0$. We have that $a=j$ and $b=j$ if for all $i\neq j$,
$\frac{-\ln(u_{j})}{p_{j}}$ $<\frac{-\ln(u_{i})}{p_{i}}$ and $\frac{-\ln(u_{j})}{q_{j}}$ $<\frac{-\ln(u_{i})}{q_{i}}.$
Equivalently, $a=j$ and $b=j$ if
$-\ln(u_{j})$ $<\min_{i\neq j}\left[-\ln(u_{i})\cdot\frac{1}{\max(p_{i}/p_{j},q_{i}/q_{j})}\right].$
Observe that $\ln(u_{j})$ is an exponential random variable with parameter $\lambda=1$ and, applying 6, $\min_{i\neq j}\left[-\ln(u_{i})\cdot\frac{1}{\max(p_{i}/p_{j},q_{i}/q_{j})}\right]$ is exponential with parameter $\lambda=\sum_{i\neq j}\max(p_{i}/p_{j},q_{i}/q_{j})$. These two random variables are independent, so again by 6:
$\Pr[a=j\text{ and }b=j]=\frac{1}{1+\sum_{i\neq j}\max(p_{i}/p_{j},q_{i}/q_{j})}.$ (7)
Summing over all $j$ for which $\min(p_{j},q_{j})>0$ yields the exactly expression for $\Pr[a=b]$ in theorem 3.
We next simplify this expression to prove that $\Pr[a=b]\geq\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}$, which establishes that Gumbel sampling at least matches the worst-case bound of Theorem 2. Assume for now that $p_{j}\leq q_{j}$. We can make a symmetric argument for the case when $q_{j}>p_{j}$. If $p_{j}\leq q_{j}$, we have:
$\Pr[a=j\text{ and }b=j]=\frac{p_{j}}{p_{j}+\sum_{i\neq j}\max(p_{i},q_{i}\cdot\frac{p_{j}}{q_{j}})}=\frac{p_{j}}{\sum_{i=1}^{n}\max(p_{i},q_{i}\cdot\frac{p_{j}}{q_{j}})}\geq\frac{p_{j}}{\sum_{i=1}^{n}\max(p_{i},q_{i})}.$
To obtain the inequality, we use that $p_{j}/q_{j}\leq 1$ when $p_{j}\leq q_{j}$. Similarly, when $q_{j}<p_{j}$, we have that $\Pr[a=j\text{ and }b=j]>\frac{q_{j}}{\sum_{i=1}^{n}\max(p_{i},q_{i})}.$ Summing over all $j$,
$\Pr[a=b]\geq\frac{\sum_{j=1}^{n}\min(p_{j},q_{j})}{\sum_{j=1}^{n}\max(p_{i},q_{i})}=\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}.$
Pareto Improvement over Weighted MinHash
Next, we strengthen the above claim by showing that, in fact, $\Pr[a=b]\geq C_{\text{WMH}}(\mathcal{P},\mathcal{Q})$. I.e., that Gumbel sampling offers a Pareto improvement over Weighted MinHash.
To prove the stronger bound, we will analyze $\Pr[a=j\text{ and }b=j]$ separately for each $j\in 1,\ldots,n$. Again, we restrict to the case when $\min(p_{j},q_{j})>0$, since if either probability equals 0, then $\Pr[a=j\text{ and }b=j]$ equals 0 for both Gumbel sampling and Weighted MinHash. When $a$ and $b$ are sampled from $\mathcal{P}$ and $\mathcal{Q}$ via Weighted MinHash, it can be seen from the analysis in Section 3 that $\Pr[a=j\text{ and }b=j]={\min(p_{j},q_{j})\left(1+|p_{j}-q_{j}|\right)}/{\sum_{i=1}^{n}\max(p_{i},q_{i})}.$ So, based on our exact expression for the collision probability for Gumbel sampling from (7), it suffices to prove that:
$\frac{\min(p_{j},q_{j})\left(1+|p_{j}-q_{j}|\right)}{\sum_{i=1}^{n}\max(p_{i},q_{i})}\leq\frac{1}{\sum_{i=1}^{n}\max\left({p_{i}}/{p_{j}},{q_{i}}/{q_{j}}\right)}.$ (8)
Summing over all $j$ would then establish our desired bound.
To prove (8), we first note that, without loss of generality, we can assume $p_{j}\leq q_{j}$ and let $q_{j}=k\cdot p_{j}$ for some $k\geq 1$. Substituting and rearranging, we can then restate the (8) as follows:
$p_{j}\left(1+(k-1)p_{j}\right)\cdot\sum_{i=1}^{n}\max\left({p_{i}}/{p_{j}},{q_{i}}/{kp_{j}}\right)$ $\leq\sum_{i=1}^{n}\max(p_{i},q_{i}).$
To prove the above, it suffices to show that:
$\sum_{i=1}^{n}S_{i}$ $\geq 0,$ where $S_{i}$ $\mathbin{\stackrel{{\scriptstyle\rm def}}{{=}}}\left(\max(p_{i},q_{i})-\left(1+(k-1)p_{j}\right)\cdot\max\left(p_{i},{q_{i}}/{k}\right)\right).$ (9)
We analyze each term in the summation by splitting into two cases:
Case 1: $p_{i}>\frac{q_{i}}{k}$.
In this case, we have:
$S_{i}=\max(p_{i},q_{i})-\left(1+(k-1)p_{j}\right)p_{i}\geq p_{i}-\left(1+(k-1)p_{j}\right)p_{i}=(1-k)p_{j}p_{i}.$
Case 2: $p_{i}\leq\frac{q_{i}}{k}$.
In this case, we have:
$S_{i}=\left(1-\frac{1}{k}-(k-1)\frac{p_{j}}{k}\right)q_{i}=\frac{k-1}{k}(1-p_{j})q_{i}\geq(k-1)(1-p_{j})p_{i}.$
We conclude that:
$\sum_{i=1}^{n}S_{i}\geq(1-k)p_{j}\cdot\left[\sum_{i:p_{i}>\frac{q_{i}}{k}}p_{i}\right]+(k-1)(1-p_{j})\cdot\left[\sum_{i:p_{i}\leq\frac{q_{i}}{k}}p_{i}\right].$
Finally, recall that, by definition, $q_{j}=kp_{j}$. So the summation $\sum_{i:p_{i}>\frac{q_{i}}{k}}p_{i}$ excludes the case $i=j$, and thus must be $\leq 1-p_{j}$. Similarly, $\sum_{i:p_{i}\leq\frac{q_{i}}{k}}p_{i}\geq p_{j}$. Recalling that $k\geq 1$, so (1k)(1-k) is non-positive and (k1)(k-1) is non-negative, we conclude that:
$\sum_{i=1}^{n}S_{i}\geq(1-k)p_{j}(1-p_{j})+(k-1)(1-p_{j})p_{j}=0.$
We have thus proven equation (9) and (8). As discussed, summing (8) over all $j$ with $\min(p_{j},q_{j})>0$ establishes that, if $a,b$ are sampled from $\mathcal{P},\mathcal{Q}$ via Gumbel sampling, then $\Pr[a=b]\geq C_{\text{WMH}}(\mathcal{P},\mathcal{Q})$. ∎
Remark.
We conclude by observing that the main inequality $\Pr[a=b]\geq C_{\text{WMH}}(\mathcal{P},\mathcal{Q})$ in Theorem 3 is actually strict under very mild conditions, i.e., in most cases Gumbel sampling strictly improves on Weighted MinHash. In particular, consider the case when there are just three indices $j,k,l$ for which $p_{j}\neq q_{j}$, and $p_{k}\neq q_{k}$, and $p_{l}\neq q_{l}$, and one of the pairs $p_{l},q_{l}$, $p_{j},q_{j}$, or $p_{k},q_{k}$ is non-zero. Then, it must be that, for two indices (without loss of generality, $j$ and $k$, either $p_{j}>q_{j}$ and $p_{k}>q_{k}$, or $q_{j}>p_{j}$ and $q_{k}>p_{k}$. Without loss of generality, assume we are in the first case. Furthermore, without loss of generality, assume that $\frac{p_{k}}{q_{k}}\geq\frac{p_{j}}{q_{j}}$.
Now, consider our proof above where we analyze $\Pr[a=j\text{ and }b=j]$. We claim that at least one of the following inequalities from the end of the proof must be strict:
$\sum_{i:p_{i}>\frac{q_{i}}{k}}p_{i}$ $\leq 1-p_{j}$ or $\sum_{i:p_{i}\leq\frac{q_{i}}{k}}p_{i}$ $\geq p_{j}.$
Specifically, if $\frac{p_{j}}{q_{j}}=\frac{p_{k}}{q_{k}}$, then $\sum_{i:p_{i}\leq\frac{q_{i}}{k}}p_{i}\geq p_{j}+p_{k}>p_{k}$, so the second inequality is strict. On the other hand, if $\frac{p_{k}}{q_{k}}>\frac{p_{j}}{q_{j}}$ then the first summation excludes both $p_{j},p_{k}$, which means $\sum_{i:p_{i}>\frac{q_{i}}{k}}p_{i}\leq 1-p_{j}-p_{k}<1-p_{j}$.
Another example discussed in Bavarian et al. (2020) is when $\mathcal{P}$ and $\mathcal{Q}$ are both uniform, but over different sized subsets (call them $\mathcal{A}$ and $\mathcal{B}$) of $\{1,\ldots,n\}$. In this special case, it can be checked that Gumbel sampling is exactly equivalent to Broder’s “unweighted MinHash” method (Broder, 1997). It yields collision probability $\frac{|\mathcal{A}\cap\mathcal{B}|}{|\mathcal{A}\cap\mathcal{B}|}$, whereas Weighted MinHash achieves a strictly worse probability of $\frac{\left(1+\left|1/|\mathcal{A}|-1/|\mathcal{B}|\right|\right)\cdot|\mathcal{A}\cap\mathcal{B}|}{|\mathcal{A}\cap\mathcal{B}|+|\mathcal{B}|-|\mathcal{A}|}$.
5 Application: Drafter-Invariant Speculative Decoding
With our main theoretical results in place, in this section we describe an application of communication-free coupling to accelerating autoregression language models, which have demonstrated impressive capabilities across a wide range of language-related tasks (Touvron et al., 2023; OpenAI, 2023; Gemini Team, Google, 2023). Given a text query $q$ (e.g., a question), the goal of autoregressive language models (LMs) is to generate a sequence of tokens $t_{1},t_{2},\ldots$, which correspond to words or word-pieces that comprise an answer or response to the query. At step $i$ of the generation, the LM computes the conditional distribution $\text{Pr}(\cdot|q,t_{1},\ldots,t_{i-1})$ and the next token is obtained by sampling according to this probability distribution. Computing this next-token distribution requires a forward pass through the neural network at every step $i$, which has significant computational costs, especially for large models.
Speculative decoding was proposed by Leviathan et al. (2023) and Chen et al. (2023) as a way to accelerate token generation. The method seeks to partially parallelize the process by using an inexpensive approximate distribution (computed using a smaller neural network) to predict, or “draft”, the next $\gamma$ tokens, $t_{i},t_{i+1},\ldots,t_{t+\gamma}$. Let the draft tokens be denoted by $\tilde{t}_{i},\tilde{t}_{i+1},\ldots,\tilde{t}_{t+\gamma}$. The larger neural network (or multiple copies of the network) can compute and sample from the distributions $\text{Pr}(\cdot|q,t_{1},\ldots,{t}_{i-1})$, $\text{Pr}(\cdot|q,t_{1},\ldots,{t}_{i-1},\tilde{t}_{i})$, $\text{Pr}(\cdot|q,t_{1},\ldots,{t}_{i-1},\tilde{t}_{i},\tilde{t}_{i+1}),\ldots,\text{Pr}(\cdot|q,t_{1},\ldots,{t}_{i-1},\tilde{t}_{i},\ldots,\tilde{t}_{i+\gamma})$ in parallel. If $\tilde{t}_{i}$ matches the token $t_{i}$ sampled from $\text{Pr}(\cdot|q,t_{1},\ldots,{t}_{i-1})$, then the sample from $\text{Pr}(\cdot|q,t_{1},\ldots,{t}_{i-1},\tilde{t}_{i})$ is a proper sample from the desired large model distribution. Further, if $\tilde{t}_{i+1}$ matches the token $t_{i+1}$ sampled from $\text{Pr}(\cdot|q,t_{1},\ldots,{t}_{i-1},\tilde{t}_{i})$, then, likewise, the sample from $\text{Pr}(\cdot|q,t_{1},\ldots,{t}_{i-1},\tilde{t}_{i},\tilde{t}_{i+1})$ is a valid sample from the large model. Overall, we can obtain $k$ samples in parallel if the first $k$ draft tokens match tokens sampled by the large model. We refer readers to Leviathan et al. (2023) for further details of the method.
In this setting, Bob corresponds to the large, expensive neural network, which knows the “true” distribution $\mathcal{Q}$ and Alice corresponds to the inexpensive neural network, which knows some distribution $\mathcal{P}$ that approximates $\mathcal{Q}$. We want Alice to sample a token from $\mathcal{P}$ that, with high probability, matches Bob’s token sampled from $\mathcal{Q}$. Alice has to generate her sample before Bob, so the setting inherently does not allow communication from Bob to Alice. However, Alice can in principal communicate information to Bob that can be used when generating his sample. This is exactly what is done in current implementations of speculative decoding: Alice actually communicates the entire distribution she used to sample $\tilde{t}_{i+j}$, which allows Bob to sample in an optimally coupled way (Leviathan et al., 2023, Algorithm 1).
The end result is that Alice’s predictions are accurate with probability exactly equal to $1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$. However, this comes with a potential caveat: the output of the process depends on both $\mathcal{P}$ and $\mathcal{Q}$. Hence, if the small “drafter” model changes, while the token distribution never changes, the exact token sampled by the large model very well could. This is undesirable: as discussed in Section 1, the output of an LM is ideally fixed given a fixed random seed, no matter what optimizations are used to accelerate inference of the network. This property allows for users to easily reproduce results and facilitates easier unit testing and debugging. Naively, speculative sampling destroys this property.
Fortunately, this issue can be fixed with a communication-free protocol! If we use a communication-free protocol to sample from the large model, then necessarily the models output is independent of the drafter model. We call the resulting approach “Drafter-Invariant Speculative Decoding”. There is a price to pay for drafter-invariance: by using a communication-free protocol, we can only ensure that the draft tokens are correct with probability $c=\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}$ instead of $c=1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$. Since the expected number of sequential tokens drafted correctly is $\frac{1}{1-c}$, this could lead to a roughly $2\times$ reduction, although as we will see, the communication-free protocols analyzed in this work, especially our Gumbel sampling approach, tend to outperform the worst-case bound of Theorem 2 for real-world distributions arising from language models.
Speculative Decoding
Drafter-Dependent (standard) Drafter-Invariant (ours)
No Drafter Drafter: Gemma 9B Drafter: Gemma 2B Drafter: Gemma 9B Drafter: Gemma 2B
Yes
Yes
Typically
Yes
Yes
It
Yes
No
some
often
often
NP-hard
NP-hard
some
but
some
NP
NP-hard
within
with
problems
problems
NP-hard
NP-hard
NP-hard
NP-hard
problems
a
a
often
can
problems
problems
problems
problems
have
bounded
prov
have
have
have
can
have
can
efficient
factor
factor
efficient
efficient
efficient
efficient
efficient
fast
approximation
of
of
approximation
approximation
approximation
approximation
approximation
approximations
algorithms.
the
the
algorithms.
algorithms.
algorithms.
algorithms.
algorithms.
algorithms.
<eos>
optimal
optimal
<eos>
<eos>
<eos>
<eos>
<eos>
<eos>
solution.
solution.
Table 1: We applied standard Speculative Decoding (Leviathan et al., 2023) and our Drafter-Invariant Speculative Decoding method to a 27 billion parameter Gemma model (Gemma Team, Google, 2024) to generate responses to the query: “Can NP-hard problems be approximated efficiently?”. Smaller 9 and 2 billion parameter Gemma models were used as drafters. The table show the tokens returned by the large model, as well as the draft tokens proposed by the smaller models (displayed below). If the draft token matches the large model, it is shown in blue; otherwise, it is show in red. As we can see, Drafter-Invariant Speculative Decoding always results in a response of “Yes some NP-hard problems have efficient approximation algorithms.”, no matter what drafter is used. In contrast, standard Speculative Decoding leads to three different responses depending on the drafter.
5.1 Experimental Evaluation
We used Gumbel sampling to implement Drafter-Invariant Speculative Decoding to accelerate generation of tokens for the 27 billion parameter gemma-2-27b-it model, using the smaller gemma-2-9b-it, and gemma-2-2b-it models as drafters (Gemma Team, Google, 2024) [^3] https://huggingface.co/docs/transformers/en/model_doc/gemma2. Table 1 shows that, for Drafter-Invariant Speculative Decoding, the tokens generated by the large model remain the same regardless of the drafter model. In contrast, this is not the case for standard Speculative Decoding, which uses an optimal coupling.
Figure 2: This plot illustrates the effectiveness of Gumbel and Weighted MinHash sampling for coupling samples. For each subplot, we obtain 32 pairs of distributions $(\mathcal{P},\mathcal{Q})$ by generating 32 response tokens for a given query using two different language models: the 27 billion parameter Gemma 27B model and the smaller Gemma 2B model. We sample $a\sim\mathcal{P}$ and $b\sim\mathcal{Q}$ 20,000 times using the no-communication protocols, and plot the empirical probability that $a=b$ vs. the total variation distance between $\mathcal{P}$ and $\mathcal{Q}$. For reference, we also plot the best possible probability that $a=b$, $1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$, as well as our bound from Theorem 2.
In Figure 2, we also experimentally compare the effectiveness of Gumbel sampling and Weighted MinHash sampling for generating coordinated samples. We issue queries to the gemma-2-27b-it model, then generate a sequence of 32 response tokens using both the full model and the gemma-2-2b-it drafter model. For each token, we have a pair of distributions, $\mathcal{Q}$ (the large model’s target distribution) and $\mathcal{P}$ (the drafter’s approximate distribution). We generate samples $a\sim\mathcal{P}$ and $b\sim\mathcal{Q}$ from these distributions using the communication-free coupling protocols. We report the fraction of times that $a=b$ over $20,000$ repetitions using different random seeds. This results in 32 plotted points, per query, per method. Since the accuracy of the drafter model varies depending on where we are in the token generation sequence, we can see that the points corresponds to distributions with a fairly wide range of total variation distances.
As we can see in the figure, Gumbel sampling achieves a higher collision probability than Weighted MinHash, across all queries and tokens, which is consistent with our Theorem 3. Both methods typically outperform the worst-case bound of $\frac{1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}{1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}$ proven in Theorem 2, often nearly matching the best possible collision probability of $1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$.
6 Coupling with Low-Communication
We conclude by considering a relaxed version of the communication-free coupling problem that allows for limited communication. In particular, Theorem 2 proves that there is a gap between communication-free protocols and the communication-intensive protocol from Protocol 1, where Alice sends her entire distribution to Bob. It is natural to ask what lies in between. Concretely, how many bits of communication are needed to match the optimal $1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$ collision probability of Protocol 1?
Our main result on this question is Theorem 4, which we restated below: See 4
To prove Theorem 4, we first introduce an “idealized” communication protocol, Protocol 4, that proceeds in rounds. At each round, Alice and Bob communicate at most one index consisting of $O(\log n)$ bits, and one probability, i.e., a real-valued number in the interval $[0,1]$. We show that this protocol terminates with an expected constant number of rounds. We then show how to appropriately discretize Alice and Bob’s distributions in a pre-processing step so that the probability can be communicated with $O(\log(n/\epsilon))$ bits. The same discretization strategy can be used to give a concrete communication complexity bound for Protocol 1, which naively requires communicating real-valued probabilities. The result would be a protocol requiring $O(n\log(n/\epsilon))$ bits of communication. Theorem 4 offers and exponential improvement on this baseline.
Protocol 4 Low Communication Coupling
Fix public random numbers $u_{0},u_{1},\ldots$ drawn uniformly from the interval $[0,1]$.
Protocol for Alice (who has probability vector $\mathcal{P}=[p_{1},\ldots,p_{n}]$):
1:Sample $a\sim\mathcal{P}$ and send $(a,p_{a})$ to Bob.
2:Await Bob’s response.
3:if Bob sends reject then
4:  for $k=1,2,\ldots$ do
5:   Await $(j,w)$ from Bob.
6:   If $p_{j}\geq u_{k}-w$, send reject to Bob. Otherwise send approve and break the loop.
7:Return $a$.
Protocol for Bob (who has probability vector $\mathcal{Q}=[q_{1},\ldots,q_{n}]$):
1:Receive $(a,p_{a})$ from Alice.
2:if $u_{0}\leq\min\left(\frac{q_{a}}{p_{a}},1\right)$ then
3:  Send approve to Alice.
4:  Return $b=a$.
5:else
6:  Send reject to Alice.
7:  for $k=1,2,\ldots$ do
8:   Find $j$ such that $\sum_{t=1}^{j-1}q_{t}\leq u_{k}<\sum_{t=1}^{j}q_{t}$ and send $(j,\sum_{t=1}^{j-1}q_{t})$ to Alice.
9:   Await Alice’s response.
10:   if Alice sends approve then
11:     Return $b=j$
Idealized Protocol. We begin by analyzing our idealized protocol, Protocol 4. It is based on a modification of the optimal coupling method, Protocol 1. While Protocol 1 naively requires communicating Alice’s entire distribution, $\mathcal{P}$, to Bob, we can see that $\mathcal{P}$ only becomes necessary if Bob “rejects” the item $a$ that he receives from Alice. In particular, upon rejectionm he must sample from a distribution $\mathcal{Q}^{\prime}$ that depends on $\mathcal{P}$.
Figure 3: Two probability distributions, $\mathcal{P}$ (red) and $\mathcal{Q}$ (blue), are depicted in this diagram. The area where $q_{i}>p_{i}$ is highlighted with a hatch pattern. In the first round of Protocol 4, Alice samples an item proportional to $\mathcal{P}$. Using rejection sampling, Bob decides whether to accept or reject the item. If rejected, Bob attempts to draw a random number that “hits” the hatched regions where $q_{i}$ exceeds $p_{i}$.
Protocol 4 implements this “post-rejection” step in a more communication efficient way by using a dart-throwing approach similar to Weighted MinHash. Specifically, Bob attempts to sample from regions of the number line where $q_{j}>p_{j}$, as illustrated by the hatched regions in Figure 3. A uniform sample from these regions amounts to a sample from $\mathcal{Q}^{\prime}$. To obtain such a sample, Bob throws a dart and, if it lands in the region $j-1,j-1+q_{j}]$ for some $j$, he sends the dart to Alice to verify that it also lies above $j-1+p_{j}$. If it does, then the dart has landed in the hatched region, so they terminate the process. If not, Bob tries again. We prove that the number of rounds required for the process to terminate is constant in expectation:
Claim 7.
If Alice and Bob use Protocol 4 to produce samples $a\sim\mathcal{P}$ and $b\sim\mathcal{Q}$, then:
$\Pr[a=b]=1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q}).$
Furthermore, the expected number of messages before the protocol terminates is at most $2$.
Proof.
Let $a$ and $b$ be the indices returned by Alice and Bob at the end of the protocol. We first prove that $a\sim\mathcal{P}$, $b\sim\mathcal{Q}$, and $\Pr[a=b]=1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$. The fact that $a\sim\mathcal{P}$ is immediate, since Alice starts by choosing $a\sim\mathcal{P}$ and then never changes $a$’s value. Alice’s selection is communicated to Bob, who then decides to either accept or reject $a$. Bob accepts with probability $\min\left(\frac{q_{a}}{p_{a}},1\right)$. So, fixing an index $i$, we have:
$\Pr[a=b=i]=p_{i}\cdot\min\left(\frac{q_{i}}{p_{i}},1\right)=\min(q_{i},p_{i}).$
Summing over all $i$ we conclude that
$\Pr[a=b]=\sum_{i=1}^{n}\Pr[a=b=i]=\sum_{i=1}^{n}\min(p_{i},q_{i})=1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q}).$
Now, if Bob rejects $a$ at Line 6, he must select another value $j\neq a$. The probability of selecting $j$ given the rejection of $a$ is:
$\Pr[b=j\mid\text{Bob rejects $a$}]=\frac{q_{j}-\min(p_{j},q_{j})}{\sum_{k=1}^{n}q_{k}-\min(p_{k},q_{k})}.$
The bottom of the fraction above is exactly the probability that Bob rejects at Line 6, so overall, $\Pr[b=j]=\Pr[b=j\land b=a]+\Pr[b=j\land b\neq a]=\min(q_{j},p_{j})+q_{j}-\min(q_{j},p_{j})=q_{j}$, as desired.
Next, we bound the round complexity of Protocol 4. Bob’s process involves a dart-throwing mechanism where each throw is aimed at hitting a region representative of the excess probability mass of $q_{j}$ over $p_{j}$. The chance of hitting the correct region per trial, shown as the hatched area in the Figure 3, is equal to the total variation distance, $D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$.
$\sum_{j=1}^{n}q_{j}-\min(p_{j},q_{j})=\sum_{j=1}^{n}q_{j}-\sum_{j=1}^{n}\min(p_{j},q_{j})=1-(1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q}))=D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q}).$
Each throw is independent, and the expected number of throws needed to hit the target region is geometrically distributed with a success probability of $D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$. Therefore, the expected number of throws needed is $\frac{1}{D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}$. Bob’s dart throwing procedure is only used if $a$ is rejected at Line 6 of Bob’s protocol, which happens with probability $D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})$. So, the expected number of communication rounds between Alice and Bob, denoted by $\mathbb{E}[c]$, is:
$\mathbb{E}[c]=\Pr[a=b]\cdot 1+\Pr[a\neq b]\cdot\frac{1}{D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}\leq 1+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})\cdot\frac{1}{D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})}\leq 2.$
With 7 in place, we are almost ready to prove Theorem 4. The only issue is that Protocol 4 must communicate probabilities, which could be arbitrarily precise real numbers, and thus require an unbounded number of bits to represent. To avoid this issue, we need to discretize $\mathcal{P}$ and $\mathcal{Q}$ using Algorithm 1, to form distributions $\mathcal{P}^{\prime}$ and $\mathcal{Q}^{\prime}$, where each probability is a multiple of $\frac{\epsilon}{n}$ for a given small $\epsilon$. This quantization allows us to communicate the probabilities using just $O(\log(n/\epsilon))$ bits.
Algorithm 1 Distribution Discretization
1:Original distribution $\mathcal{P}=\{p_{1},\ldots,p_{n}\}$, precision parameter $\epsilon$, and size $n$.
2:Distribution $\mathcal{P}^{\prime}=\{p_{1}^{\prime},\ldots,p_{n}^{\prime}\}$ where each $p_{i}^{\prime}$ is a multiple of $\epsilon/n$.
3:for $p_{i}\in\mathcal{P}$ do
4:  Set $p_{i}^{\prime}$ equal to $p_{i}$ rounded down to the nearest multiple of $\frac{\epsilon}{n}$.
5:Set $r\leftarrow(1-\sum_{i=1}^{n}p_{i}^{\prime})$.
6:Set $p_{1}^{\prime}\leftarrow p_{1}^{\prime}+r$. $\triangleright$ This step ensures that $\sum_{i=1}^{n}p_{i}^{\prime}=1$.
Lemma 8.
For any distributions $\mathcal{P}$ and $\mathcal{Q}$ over $\{1,\ldots,n\}$, if we apply Algorithm 1 with parameters $\epsilon,n$ to get distributions $\mathcal{P}^{\prime}$ and $\mathcal{Q}^{\prime}$, then we have:
$D_{\mathrm{TV}}(\mathcal{P},\mathcal{P}^{\prime})$ $\leq\epsilon,$ $D_{\mathrm{TV}}(\mathcal{Q},\mathcal{Q}^{\prime})\leq\epsilon,$ and $D_{\mathrm{TV}}(\mathcal{P}^{\prime},\mathcal{Q}^{\prime})$ $\leq D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})+2\epsilon.$
Proof.
First, consider $\mathcal{P}$. Let $r$ be as in Line 3 of Algorithm 1. We have:
$r=1-\sum_{i=1}^{n}p_{i}^{\prime}\leq 1-\sum_{i=1}^{n}\left(\frac{p_{i}}{\frac{\epsilon}{n}}-1\right)\cdot\frac{\epsilon}{n}=1-\sum_{i=1}^{n}\left(p_{i}-\frac{\epsilon}{n}\right)=\sum_{i=1}^{n}\frac{\epsilon}{n}=\epsilon.$
We can then bound the total variation distance between $\mathcal{P}$ and $\mathcal{P}^{\prime}$ as:
$D_{\mathrm{TV}}(\mathcal{P},\mathcal{P}^{\prime})$ $=\frac{1}{2}\sum_{i=1}^{n}|p_{i}-p^{\prime}_{i}|\leq\frac{1}{2}\cdot 2r\leq\epsilon.$
We can similarly demonstrate that $D_{\mathrm{TV}}(\mathcal{Q},\mathcal{Q}^{\prime})<\epsilon$. Finally, the last bound follows from triangle inequality:
$D_{\mathrm{TV}}(\mathcal{P}^{\prime},\mathcal{Q}^{\prime})$ $\leq D_{\mathrm{TV}}(\mathcal{P}^{\prime},\mathcal{P})+D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})+D_{\mathrm{TV}}(\mathcal{Q},\mathcal{Q}^{\prime})\leq D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})+2\epsilon$
Our final approach will be for Alice and Bob to run Protocol 4 on the distributions $\mathcal{P}^{\prime}$ and $\mathcal{Q}^{\prime}$ instead of $\mathcal{P}$ and $\mathcal{Q}$. This produces samples $a\sim\mathcal{P}^{\prime},b\sim\mathcal{Q}^{\prime}$ instead of $a\sim\mathcal{P},b\sim\mathcal{Q}$, as required. To correct the distributions of the samples, Alice and Bob simply use an optimal coupling between $\mathcal{P}$ and $\mathcal{P}^{\prime}$ and between $\mathcal{Q}$ and $\mathcal{Q}^{\prime}$. These couplings can be computed locally, without any communication.
Proof of Theorem 4.
Alice and Bob round their distributions using Algorithm 1 with parameter $\epsilon/4$ to obtain $\mathcal{P}^{\prime}$ and $\mathcal{Q}^{\prime}$. They then run Protocol 4 with $\mathcal{P}^{\prime}$ and $\mathcal{Q}^{\prime}$ to obtain $a^{\prime}\sim\mathcal{P}^{\prime}$ and $b^{\prime}\sim\mathcal{Q}^{\prime}$ such that $\Pr[a^{\prime}=b^{\prime}]\geq 1-D_{\mathrm{TV}}(\mathcal{P}^{\prime},\mathcal{Q}^{\prime})$. They each separately apply Protocol 1 to obtain samples $a\sim\mathcal{P}$ and $b\sim\mathcal{Q}$ such that $\Pr[a=a^{\prime}]\geq 1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{P}^{\prime})$ and $\Pr[b=b^{\prime}]\geq 1-D_{\mathrm{TV}}(\mathcal{Q},\mathcal{Q}^{\prime})$. Overall, we have:
$\Pr[a=b]$ $\geq\Pr[a=a^{\prime}\text{ and }a^{\prime}=b^{\prime}\text{ and }b=b^{\prime}]$
$\geq 1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{P}^{\prime})-D_{\mathrm{TV}}(\mathcal{Q}^{\prime},\mathcal{P}^{\prime})-D_{\mathrm{TV}}(\mathcal{Q},\mathcal{Q}^{\prime})$ (by a union bound)
$\geq 1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})-4\cdot\epsilon/4$ (by Lemma 8)
$=1-D_{\mathrm{TV}}(\mathcal{P},\mathcal{Q})-\epsilon.$
The total communication cost is that of running Protocol 4 with $\mathcal{P}^{\prime}$ and $\mathcal{Q}^{\prime}$. Each round of this protocol requires communicating an index, which takes $O(\log n)$ bits, and a probability, which takes $O(\log(n/\epsilon))$ bits (since all probabilities are integer multiples of $\epsilon/4n$). By 7, the protocol terminates after a constant number of rounds in expectation, so in total, we need $O(\log(n/\epsilon))$ bits of communication, in expectation. ∎
Acknowledgement
This work was partially support by NSF Grant #2106888\#2106888. Authors thank Mert Cemri, Nived Rajaraman, and Cheuk Ting Li for pointing out related works.